Comet d'Arrest was discovered in June 28, 1851 and its orbit about the sun has a semimajor axis a = 3.49 AU. Recently it was observed that the comet has a velocity vc = 0.45V E ev when it was spotted at position rc = 2.91RE er from the sun at which the velocity vector vc makes an angle ϕ = 60o with vector rc (VE and RE = 1 AU are the average speed and radius of the Earth orbital motion around the sun respectively). Calculate i. the eccentricity of the comet's orbit. ii. the period of the comet in Earth years. iii. the perihelion, aphelion and the latus rectum of the orbit. 


1
Expert's answer
2021-02-22T10:21:37-0500

1)

"e=\frac{r_c}{a}sin^2 \phi \approx 0.625,"

2)

"T=a^{1.5}\approx 6.520 ~a,"

3)

"r_p=a(1-e)=1.309 ~AU,"

"r_a=a(1+e)=5.671~AU,"

"p=a(1-e^2)=2.127~AU."